OpenStudy (anonymous):

Find value of tan theta:

OpenStudy (anonymous):

\[35+\sec^2\theta = 12\tan \theta\]

OpenStudy (shubhamsrg):

Use the identity : sec^2 x = 1 + tan^2 x

OpenStudy (anonymous):

\[35 + \sec^ \theta = 12 \tan \theta \] \[36+\tan^2 \theta = 12 \tan \theta\] Not sure how to proceed, am pretty new to this and my finals are next week :(

OpenStudy (anonymous):

put \(tan\ \theta = x\) You have \(x^2 - 12x +36 = 0\) Solve this equation for x and hence for \(tan\ \theta\).

OpenStudy (anonymous):

Got it, thanks!